OpenStudy (anonymous):

use factoring to find the zeros of y=x^2+2x-24

OpenStudy (turingtest):

x={-6,4}

OpenStudy (anonymous):

find two numbers that multiply to be -24 and add to be 2

OpenStudy (anonymous):

well, there you go

OpenStudy (turingtest):

i felt like they didn't care for an explanation

OpenStudy (anonymous):

thanks having trouble understanding my work

OpenStudy (turingtest):

x^2+2x-24=0 (x-4)(x+6)=0 can you not factor?

OpenStudy (anonymous):

the best way to answer is to use the x and y plot or factorize it.